avasargent25:

Solve the following system of equations: −2x + y = 1 −4x + y = −1

D4nniii:

i am no good for this

D4nniii:

its 1,3

snowflake0531:

(since the user went offline) you can make both y equals something \( y=2x+1\) and \(y=4x-1\) so then you can use substitution and get \(2x+1 = 4x-1 \\ 2x=2\\x=1\) then replace x and get y \(y=2(1)+1\\y=2+1 \\y=3\) so you get (1,3)

D4nniii:

well ur good at math. good job i could never